A cheering squad is made up of three members: Lynne of 50kg moving at 3 m/sNorth. Angie of mass 58kg moving eastward at 3 m/s and Mercy of mass 6 kg moving at 3 m/s south. What is the combined momentum of the squad?


1
Expert's answer
2021-12-21T12:15:40-0500

Find the combined momentum. The momentum has direction equal to the direction of velocity, the equation is "p=mv."


"p_L=50·3=150\text{ kg}·\text{m/s N}.\\
p_A=58·3=174\text{ kg}·\text{m/s E}.\\
p_M=60·3=180\text{ kg}·\text{m/s S}.\\
p_\text{North}=p_L-p_M=150-180=-30\text{ kg}·\text{m/s N}.\\
p_\text{East}=174\text{ kg}·\text{m/s N}.\\\space\\
p_\text{Combined}=\sqrt{p_\text{East}^2+p_\text{North}^2}=177\text{ kg}·\text{m/s N}.\\\space\\
\theta=\arctan\frac{p_\text{North}}{p_\text{East}}=9.8°\text{ S of E}."